# Gaussian Adaptive Attention is All You Need: Robust Contextual Representations Across Multiple Modalities

### 25 Jan 2024 v2

@@ 3.3 Encoder and Decoder models
We apply different Attention mechanisms on SSL-based PTMs acting as Encoders to extract Embeddings from. Specifically, we utilize the pre-trained model weights from three distinct Encoders: (i) WavLM-Large, (ii) Llama2-13B, and (iii) BEiT-Large. Libri-light [?] GigaSpeech [?] and English parts of VoxPopuli [?] have been used for pre-training (i). ImageNet-1k [?] has been used for pre-training (iii). Conversely, while (ii) has undergone pre-training on undisclosed, publicly sourced textual data, this aspect does not impact our research. The downstream application we employ is different from the model’s original pre-training task. The specific datasets used in this work are described in further detail in Section 4.1.
@@ 4.2 Current benchmarks
Method Best Acc. MHA 0.604 MHA $\rightarrow$ BN 0.630 GAAMv1 0.799 GQGAAM 0.800 SGA 0.802 bi-SGA 0.802 GAAMv2 0.802
Method Best Acc. MHA 0.944 GAAMv2 0.944 MHA $\rightarrow$ BN 0.945 SGA 0.945 bi-SGA 0.945 GAAMv1 0.945 GQGAAM 0.948
@@
In Tables 2-4, we also contrast GAAM-based methods with those using Batch Normalization (BN), Skewed Gaussian Attention (SGA) and bi-Skewed Gaussian Attention (bi-SGA) [?]. BN operates under the assumption of independent and identical distribution across mini-batches [?] and we position BN immediately after MHA to match the normalization order in the original Transformer. GAAM, with its Multi-Headed nature, adeptly handles variations in feature distribution, including shifts in mean offset, variance, and modelling a mixture of Gaussian distributions with varying attention weights as evidenced in Table 5 and Figure 2 respectively. This adaptability leads to GAAM’s superior performance over methods reliant on static feature distributions such as BN, SGA and bi-SGA. SGA and bi-SGA assume data strictly follows a single Gaussian distribution at a time, limiting its ability to model non-Gaussian traits by not having a learnable mixture of them. In contrast, GAAM can model multiple Gaussian distributions with varying parameters, enabling it to approximate any Probability Distribution.
Results show that GAAM outperforms in scenarios with significant variation in Gaussian parameters (i.e. need for modelling non-stationary data, perhaps having non-Gaussian traits overall), while in cases with minimal variation, outperforms MHA and matches (bi-)SGA, depending on the number of heads used.
Observing Table 5, Speech data exhibits high variability in both the central tendency (mean offset, $\delta$) and the spread (scaled variance, $\xi$). This reflects the highly non-stationary nature of speech [?]. Attention mechanisms must adaptively adjust both the focus ($\mu$) and the width ($\sigma$) to effectively capture the rapidly changing features in speech, essential for tasks like emotion recognition.
Text data shows high mean variation perhaps due to changing semantic contexts, but the variance remains moderately stable, reflecting the structured nature of language. Attention mechanisms should primarily focus on tracking the shifting mean ($\mu$) to align with the changing semantic and syntactic focal points, while the moderate variance variation allows for a relatively consistent attention span.
In Vision, both the mean and variance variations are low, indicating stable and consistent feature locations and spreads. Simpler attention mechanisms can be effective, maintaining a consistent focus ($\mu$) and width ($\sigma$), suitable for tasks where feature variability is minimal.
@@ 4.4 Ablation Study
Across Speech, Text, and Vision modalities, higher IF scores consistently align with improved model performance and vice versa. For the Speech downstream task, High IF layer (Layer 9) consistently outperforms low IF layer (Layer 23) across all folds (Table 6). For the Text and Vision downstream tasks (Table 7), layers with higher IF achieve better performance, especially in Vision.
In MHA, attention weights primarily indicate the level of correlation between different parts of the input sequence [?]. Each element’s weight reflects its relevance to every other element within the same sequence. However, this approach does not directly translate to the performance on downstream tasks. For instance, the authors in [?] derive normalized self-attention weights for SER on IEMOCAP using WavLM-Large, identifying layer 23 as pivotal. While useful for their use case, this only indicates inter-layer correlation, and not a direct link to better or worse performance and would be misleading to use these weights for that use case (as shown in Table 6 using GAAM for the same task). In contrast, GAAM-based learning dynamically adjusts attention weights tailoring attention to improve feature representation aligned with the model’s end goal. Analysis of Figure 3 indicates earlier layers, exhibit more meaningful features, and contribute more to model performance, suggesting potential overparameterization in later layers [?].
Future work should explore GAAM in additional tasks, datasets, grounding experiments [?], and beyond feature extraction, including model compression using attention weights during training (crucial for resource-limited applications) [?].

### 31 Jan 2024 v3

@@ 3.3 Encoder and Decoder models
We apply different Attention mechanisms on SSL-based PTMs acting as Encoders to extract Embeddings from. Specifically, we utilize the pre-trained model weights from three distinct Encoders: (i) WavLM-Large, (ii) Llama2-13B, and (iii) BEiT-Large.
Libri-light [?] GigaSpeech [?] and English parts of VoxPopuli [?] have been used for pre-training (i). ImageNet-1k [?] has been used for pre-training (iii). Conversely, while (ii) has undergone pre-training on undisclosed, publicly sourced textual data, this aspect does not impact our research. The downstream application we employ is different from the model’s original pre-training task. The specific datasets used in this work are described in further detail in Section 4.1.
@@ 4.2 Current benchmarks
Method Best Acc.
MHA 0.604
MHA $\rightarrow$ BN 0.630
GAAMv1 0.799
GQGAAM 0.800
SGA 0.802
bi-SGA 0.802
GAAMv2 0.802
Method Best Acc.
MHA 0.944
GAAMv2 0.944
MHA $\rightarrow$ BN 0.945
SGA 0.945
bi-SGA 0.945
GAAMv1 0.945
GQGAAM 0.948
@@
In Tables 2-4, we also contrast GAAM-based methods with those using Batch Normalization (BN), Skewed Gaussian Attention (SGA) and bi-Skewed Gaussian Attention (bi-SGA) [?]. BN operates under the assumption of independent and identical distribution across mini-batches [?] and we position BN immediately after MHA to match the normalization order in the original Transformer. GAAM, with its Multi-Headed nature, adeptly handles variations in feature distribution, including shifts in mean offset, variance, and modelling a mixture of Gaussian distributions with varying attention weights as evidenced in Table 5 and Figure 2 respectively. This adaptability leads to GAAM’s superior performance over methods reliant on static feature distributions such as BN, SGA and bi-SGA. SGA and bi-SGA assume data strictly follows a single Gaussian distribution at a time, limiting its ability to model non-Gaussian traits by not having a learnable mixture of them. In contrast, GAAM can model multiple Gaussian distributions with varying parameters, enabling it to approximate any Probability Distribution. Results show that GAAM outperforms in scenarios with significant variation in Gaussian parameters (i.e. need for modelling non-stationary data, perhaps having non-Gaussian traits overall), while in cases with minimal variation, outperforms MHA and matches (bi-)SGA, depending on the number of heads used. Observing Table 5, Speech data exhibits high variability in both the central tendency (mean offset, $\delta$) and the spread (scaled variance, $\xi$). This reflects the highly non-stationary nature of speech [?]. Attention mechanisms must adaptively adjust both the focus ($\mu$) and the width ($\sigma$) to effectively capture the rapidly changing features in speech, essential for tasks like emotion recognition. Text data shows high mean variation perhaps due to changing semantic contexts, but the variance remains moderately stable, reflecting the structured nature of language. Attention mechanisms should primarily focus on tracking the shifting mean ($\mu$) to align with the changing semantic and syntactic focal points, while the moderate variance variation allows for a relatively consistent attention span. In Vision, both the mean and variance variations are low, indicating stable and consistent feature locations and spreads. Simpler attention mechanisms can be effective, maintaining a consistent focus ($\mu$) and width ($\sigma$), suitable for tasks where feature variability is minimal.
@@ 4.4 Ablation Study
Across Speech, Text, and Vision modalities, higher IF scores consistently align with improved model performance and vice versa. For the Speech downstream task, High IF layer (Layer 9) consistently outperforms low IF layer (Layer 23) across all folds (Table 6). For the Text and Vision downstream tasks (Table 7), layers with higher IF achieve better performance, especially in Vision. In MHA, attention weights primarily indicate the level of correlation between different parts of the input sequence [?]. Each element’s weight reflects its relevance to every other element within the same sequence. However, this approach does not directly translate to the performance on downstream tasks. For instance, the authors in [?] derive normalized self-attention weights for SER on IEMOCAP using WavLM-Large, identifying layer 23 as pivotal. While useful for their use case, this only indicates inter-layer correlation, and not a direct link to better or worse performance and would be misleading to use these weights for that use case (as shown in Table 6 using GAAM for the same task). In contrast, GAAM-based learning dynamically adjusts attention weights tailoring attention to improve feature representation aligned with the model’s end goal. Analysis of Figure 3 indicates earlier layers, exhibit more meaningful features, and contribute more to model performance, suggesting potential overparameterization in later layers [?]. Future work should explore GAAM in additional tasks, datasets, grounding experiments [?], and beyond feature extraction, including model compression using attention weights during training (crucial for resource-limited applications) [?].
@@ 5 Conclusion
6 Multi-Head Gaussian Adaptive Attention Mechanism Extension
This section presents an extension of the Multi-Head Gaussian Adaptive Attention Mechanism (GAAM), focusing on enhancing the stability of the training process and the model’s efficiency by reducing the number of learnable parameters. The proposed method integrates multi-head attention mechanisms with Gaussian mixtures and skip connections to provide a more refined and adaptable approach to handling complex datasets.
The extended GAAM incorporates multiple attention heads, each with its Gaussian mixture model, to process different segments of the input tensor in parallel. This approach allows for a more diverse and comprehensive understanding of the data, leading to increased model robustness and efficiency.
6.1 Algorithmic Details
Gaussian Adaptive Attention
This algorithm which forms the core of the extended GAAM, implementing the Gaussian mixture model within each attention head can be found in listing 3. Key elements include initialization of Gaussian parameters and mean offsets, and a forward pass handling.
Algorithm 3 Gaussian Adaptive Attention Mechanism
0:  $x$ (input tensor), normAxis, numGaussians, eps
0:  Attention-modified tensor
1:  Initialize an array meanOffsets of size numGaussians
2:  Initialize an array c of size numGaussians
3:  $\text{mean}\leftarrow\text{mean}(x,\text{axis}=\text{normAxis})$
4:  $\text{var}\leftarrow\text{var}(x,\text{axis}=\text{normAxis})+\text{eps}$
5:  $\text{mixture}\leftarrow 1$
6:  for $i=0$ to $\text{numGaussians}-1$ do
[…]
8:     $\text{yNorm}\leftarrow(\text{x}-\text{adjustedMean})/\sqrt{\text{var}}$
9:     $\text{gaussian}\leftarrow\exp(-(\text{yNorm}^{2})/(2\cdot\textbf{c}[i]^{2}))/\sqrt{2\pi\textbf{c}[i]^{2}}$
[…]
11:  end for
12:  Normalize mixture across normAxis
[…]
14:  return  x
Multi Head Gaussian Adaptive Attention
This part of the framework remains the same as before (refer to listing 2).
Gaussian Block
The Gaussian Block (listing 4) encapsulates multiple layers of Multi-Head Gaussian Adaptive Attention, each layer processing the input tensor independently integrating a skip connection between them to enhance stability during training and inference.
Algorithm 4 Gaussian Block
0:  $x$ (input tensor), normAxes, numHeads, numGaussians, paddingValue, eps
0:  Final modified tensor
1:  for each layer in MultiHeadGaussianAdaptiveAttention do
2:     $x$leftarrow layer($x$) + $x$
3:  end for
4:  return  $x$